# Cooling and heating regions of Joule-Thomson expansion for AdS black holes: Maxwell-power-Yang-Mills and Kerr Sen black holes

### v1

@@ 4 Case II: AdS-Kerr-Sen black hole
The Kerr-Sen-AdS black hole is a complex and fascinating topic in the field of theoretical physics. It is a type of rotating, charged black hole that emerges from heterotic string theory. It is a generalization of the Kerr-Newman-AdS black hole, which is a solution of the Einstein-Maxwell equations with a negative cosmological constant. The Kerr-Sen-AdS black hole also involves a dilaton and an axion field, which are scalar and pseudoscalar fields that appear in string theory [61]. The first exact solution of the Einstein field equations, known as the Schwarzschild solution, was discovered by Karl Schwarzschild in 1916. This solution describes a simple, non-rotating, uncharged black hole. The next major advancement came in 1963, when Roy P. Kerr found a solution to the Einstein field equations that describes a rotating black hole. This was a significant development, as it is believed that most black holes in the universe are rotating. In 1965, Ezra Newman and his collaborators found a solution that describes a rotating, charged black hole, known as the Kerr-Newman black hole. This solution was later extended by Ashoke Sen to include a dilaton field and an axion field, resulting in the Kerr-Sen black hole. The AdS form of this black hole was first derived by Ashoke Sen in 1992, by applying a series of duality transformations to the Kerr-Newman-AdS black hole. Sen showed that the Kerr-Sen-AdS black hole retains some of the symmetries and properties of the Kerr-Newman-AdS black hole, such as the existence of an event horizon, an ergosphere, and a Penrose process. However, the Kerr-Sen-AdS black hole also has some unique features, such as the dependence of the mass and angular momentum on the dilaton charge, and the violation of the cosmic censorship conjecture for some values of the parameters [61]. The Kerr-Sen-AdS black hole solution is a specific example of a rotating black hole with additional fields, which is of particular interest due to the role of angular momentum and the presence of nontrivial matter content in the spacetime geometry and is a solution to the equations of motion of supergravity theories with additional matter fields, such as the Sen-type dilaton and antisymmetric tensor fields, within the context of AdS space.
The properties and thermodynamics of Kerr Sen-AdS black holes have been the subject of intense research, given their relevance to understanding the behavior of rotating black holes in the presence of nontrivial matter content and their implications for the AdS/CFT correspondence[62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. The study of Kerr Sen-AdS black holes has been motivated by theoretical developments in string theory, quantum gravity, and holography, as well as by their potential implications for gravity dualities and the behavior of strongly coupled systems in the dual field theories. This black hole has been studied from various perspectives, including its thermodynamic properties, stability, and connections to the dynamics of dual field theories. In recent years, there has been growing interest in exploring the dynamical behavior of Kerr Sen-AdS black holes, including their evolution, instability, and the connections to chaos and information loss puzzles. Researchers have investigated the behavior of these black holes under various perturbations and have sought to understand their implications for fundamental questions about black hole thermodynamics and the fate of information in quantum gravity. Moreover, the holographic interpretation of Kerr Sen-AdS black holes has led to fruitful connections between gravitational physics and field theory, shedding light on fundamental questions in quantum gravity and strongly coupled systems[62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. In summary, the history and ongoing research on Kerr Sen-AdS black holes represent a rich and interdisciplinary area of study that has fruitful connections to diverse fields of theoretical physics, including string theory, quantum gravity, holography, and nonlinear dynamics. The exploration of these black holes continues to be an exciting frontier for probing the fundamental nature of spacetime and its connections to quantum theory. In this section provides a short overview of the Kerr-Sen black hole and its generalization to the anti-de Sitter spacetimes. Sen [61] found a solution of the low-energy effective action of the heterotic string theory, which describes a charged rotating black hole, known as the Kerr-Sen black hole.The action is a modification of the general relativity action with additional fields from the heterotic string theory, given by[62, 78],
@@
$\begin{split}S=\int d^{4}x\sqrt{-\widetilde{g}}e^{-\Phi}\left[\mathcal{R}+(\nabla\Phi)^{2}-\frac{1}{8}F^{2}-\frac{1}{12}H^{2}\right],\end{split}$ (9)
where $\tilde{g}$ is the determinant of the metric tensor $g_{\mu\nu}$, $\mathcal{R}$ is the Ricci scalar, $F=F_{\mu\nu}F^{\mu\nu}$ with $F_{\mu\nu}$ being the $U(1)$ Maxwell field strength tensor, $\Phi$ is a scalar dilaton field, and $H=H_{\mu\nu\rho}H^{\mu\nu\rho}$ is the field strength for the axion field. The action can be transformed to the Einstein frame by a conformal transformation of the metric:
$\begin{split}ds^{2}_{E}=e^{-\Phi}d\widetilde{s}^{2}.\end{split}$ (10)
The action in the Einstein frame is given by:
$\begin{split}S=\int d^{4}x\sqrt{-g}\left[R-\frac{1}{2}(\nabla\Phi)^{2}-\frac{e^{-\Phi}}{8}F^{2}-\frac{e^{-2\Phi}}{12}H^{2}\right].\end{split}$ (11)
The metric of a Kerr-Sen-AdS black hole in Boyer-Lindquist coordinates is given by[62, 78]:
$\begin{split}ds^{2}=-\frac{\Delta_{r}}{\rho^{2}}(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}[-\frac{(r^{2}+2br+a^{2})}{\Xi}d\phi+adt]^{2},\end{split}$ (12)
@@
$\begin{split}&\rho^{2}=r^{2}+a^{2}\cos^{2}\theta+2br,\\ &\Delta=(r^{2}+2br+a^{2})(1+\frac{r^{2}+2br}{\ell^{2}})-2mr,\\ &\Xi=1-\frac{a^{2}}{\ell^{2}},\\ &\Delta_{\theta}=1-\frac{a^{2}}{\ell^{2}}\cos^{2}\theta.\end{split}$ (13)
The parameter b is the dyonic charge of the black holes and is expressed as $b=q^{2}/(2m)$, where q is the electric charge and m is the mass of the black holes. In the limit of $\ell\to\infty$, the metric reduces to the usual Kerr-Sen black holes. The non-rotating case ($a=0$) reduces to the Gibbons-Maeda-Garfinkle-Horowitz-Strominger (GMGHS) solution. Gibbons and Maeda [73, 74] found the black hole and black brane solutions in the dilaton theory, and Garfinkle-Horowitz-Strominger obtained their charged version [75]. It is worth mentioning that the Kerr-Sen-AdS black holes in four dimensions have been explored from various aspects, such as the black hole shadows [76] and the phase space thermodynamics in the extended phase space[77]. The ADM mass M, the angular momentum J, and the charge Q in AdS spacetimes are related by,
$\begin{split}&M=\frac{m}{\Xi^{2}},\\ &J=\frac{ma}{\Xi^{2}},\\ &Q=\frac{q}{\Xi}.\end{split}$ (14)
Also, the entropy is given by,
$\begin{split}S=\frac{A}{4}=\frac{\pi(r_{+}^{2}+2br+a^{2})}{\Xi},\end{split}$ (15)
where $r_{+}$ is the radius of the event horizon, which is the largest root of $\Delta=0$. When b=0, the thermodynamic quantities become the same as those of Kerr-AdS black holes. Kerr-AdS black holes are solutions of the Einstein field equations in four dimensions with negative cosmological constant and rotation.
@@ 5 Discussion and results
In this section, we investigate the JTE process for AdS black holes: AMPYM and AdS Kerr Sen and derive a clear formula for the JTC. We also examine how the charge and the parameters of AMPYM and Kerr-Sen theories affect the inversion curves. We also compare the inversion curves for different scenarios
@@ 5.1 Case I
In this section, we explore the JTE of black hole systems in the extended phase space, where the black hole mass M is the same as the enthalpy H. We compare this process with the JT process of van der Waals gases with fixed particle number, and we use the fixed charge Q for the black hole systems. We also assume that the other parameters are constant. Using [2, 45] as a reference, and considering the mass of black holes, we can write the pressure P as a function of M and $r_{+}$, the mass and the horizon radius of a black hole. By substituting this expression in the temperature formula, we can obtain the temperature as a function of M and $r_{+}$ as well. Furthermore, we can express the mass M and the temperature T in terms of the pressure P and the radius $r_{+}$ of a black hole. Therefore, we get,
$P(M,r_{+})=-\frac{3\left(\frac{2C^{2}(n-1)r_{+}^{-2n}}{n}+Qr_{+}^{-4q}+\frac{1}{r_{+}^{2}}\right)}{8\pi}+\frac{6Mr_{+}^{-n-1}}{n\omega_{n}}.$ (16)
@@
$T(M,r_{+})=-\frac{2C^{2}(n-1)r_{+}^{2-2n}+2qQr_{+}^{2-4q}+1}{2\pi r_{+}}+\frac{4M(n+1)r_{+}^{-n}}{n\omega_{n}}.$ (17)
By solving the above equations, one can obtain the function T(M, P), which is lengthy and will not be shown here. For particular free parameters for this model, the T(M, P) curve can be shown. According to the definition of the JTC $\mu=\left(\frac{\partial T}{\partial P}\right)_{M}$, the inversion pressure and temperature between the cooling and heating regions are ($\tilde{P}$-$\tilde{T}$), which are determined by $\mu=0$. Therefore, the most important thing is to find the function expression of $\mu$. By setting $\mu=0$, one can obtain the inversion points ($\tilde{P}$, $\tilde{T}$) for different fixed enthalpy M. With respect to[24], the JTC is given by,
$\begin{split}\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{M}=\frac{1}{C_{P}}\big{[}T(\frac{\partial V}{\partial T}\big{)}_{P}-V\big{]}.\end{split}$ (18)
@@
$\begin{split}\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{M}=\frac{\partial_{,r_{+}}T}{\partial_{,r_{+}}P}.\end{split}$ (19)
@@
$\begin{split}&\mu=\frac{2nr_{+}\left(\omega_{n}\left(r_{+}^{4q}\left(2C^{2}\left(2n^{2}-3n+1\right)r_{+}^{2}+r_{+}^{2n}\right)+8q^{2}Qr^{2n+2}-2qQr_{+}^{2n+2}\right)-8\pi M(n+1)r_{+}^{n+4q+1}\right)}{3\left(n\omega_{n}\left(r_{+}^{4q}\left(2C^{2}(n-1)r_{+}^{2}+r_{+}^{2n}\right)+2qQr_{+}^{2n+2}\right)-8\pi M(n+1)r^{n+4q+1}\right)}.\end{split}$ (20)
@@
$T_{i}=V\big{(}\frac{\partial T}{\partial V}\big{)}_{P},$ (21)
$\begin{split}V=\frac{\partial M}{\partial p}=\frac{1}{6}nr_{+}^{n+1}\omega_{n}.\end{split}$ (22)
@@
$\begin{split}T_{i}=V\frac{\partial T}{\partial V}=V\frac{\frac{\partial T}{\partial r_{+}}}{\frac{\partial V}{\partial r_{+}}}.\end{split}$ (23)
@@
This is beneficial for identifying the areas of heating and cooling in the $T-P$ plane. We calculate t using the equations (17), (20), (21), (22) and (23),
$\begin{split}T_{i}=\frac{r_{+}\left(\frac{6C^{2}(n-1)^{2}(2n-1)r_{+}^{-2n}}{n}+8\pi(n+1)P_{i}+3(1-4q)Q(n-4q+1)r_{+}^{-4q}+\frac{3-3n}{r_{+}^{2}}\right)}{12\pi(n+1)}.\end{split}$ (24)
We can also have from relation (17),
$\begin{split}T_{i}=\frac{-\frac{6C^{2}(n-1)^{2}r_{+}^{1-2n}}{n}+8\pi(n+1)r_{+}P_{i}+3Q(n-4q+1)r_{+}^{1-4q}+\frac{3(n-1)}{r_{+}}}{12\pi}.\end{split}$ (25)
@@
Figure 1 displays the Hawking temperature as a function of the horizon. We keep the free parameters constant for each plot. In each subfigure, we can observe some zero points for different free parameters. These zero points correspond to the divergence points of the JTC, as we can easily see in figure 1. According to the radius of the horizon and the values of free parameters of a black hole, for small values for the radius of the event horizon, our structural behavior is completely distinct, and for larger radii, the figures almost converge.
Figure 1: It shows the plot of $(T-r_{+})$ for the AMPYM black hole with respect to free parameters that are determined in each plot
@@
We have plotted the isenthalpic curves and the inversion curve of the AMPYM black hole for various values of the free parameter in each plot of Figure 2. In every subfigure, two isenthalpic curves with different values of M, along with the corresponding inversion curve that occurs at the highest point of the isenthalpic curves. We denote the inversion temperature and pressure of each isenthalpic curve as $T_{i}$ and $P_{i}$, respectively. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole undergoes cooling during the expansion process. However, for $P>P_{i}$, the isenthalpic curve has a negative slope, implying that the black hole experiences heating during the expansion process. This behavior is consistent for different values of the free parameter in Figures (2a-2b).
Figure 2: Isenthalpic curves and inversion curve of the AMPYM black hole with respect to mentioned free parameters that determine in each plot.
@@
We do not present the explicit expressions of the solutions here, as they are too long and complex. However, we can illustrate the relation between the inversion temperature $T_{i}$ and the inversion pressure $P_{i}$ by using the equations given above. Figure 3 shows the inversion curves for different free parameters. The inversion curve has only one branch. The inversion temperature rises steadily with the inversion pressure, but the slope of the inversion curves becomes smaller. We can also notice some fine structures in the cases of subfigures 3. For low pressure, the inversion temperature varies with the free parameters that are specified in each plot. It was interesting for us to study the effect of parameters and we observed that the slope of the inversion curve increases with the changes in the various values of each of the parameters.
Figure 3: It shows the plot of The inversion curve for the AMPYM black hole with respect to free parameters that are determined in each plot

### v2

@@ 4 Case II: AdS-Kerr-Sen black hole
The KS black hole, a rotating and charged black hole from heterotic string theory, generalizes the Kerr-Newman-AdS black hole. It includes a dilaton and an axion field, which are scalar and pseudoscalar fields from string theory. The Schwarzschild solution, discovered in 1916, describes a simple black hole, while the Kerr solution in 1963 introduced rotating black holes. The Kerr-Newman black hole, found in 1965, added charge, and Ashoke Sen extended it to include dilaton and axion fields, resulting in the Kerr-Sen black hole. The AKS black hole retains some properties of the Kerr-Newman-AdS black hole but has unique features like mass and angular momentum dependence on the dilaton charge. It also sometimes violates the cosmic censorship conjecture. The KS black hole’s properties and thermodynamics have been extensively studied, especially in the context of the AdS/CFT correspondence. Research has focused on their thermodynamic properties, stability, and connections to dual-field theories. The holographic interpretation of KS black holes has linked gravitational physics with field theory, addressing fundamental questions in quantum gravity and strongly coupled systems. This ongoing research continues to explore the dynamic behavior, evolution, and implications for chaos and information loss puzzles in quantum gravity[90, 91, 92, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104].
Sen’s work [90] unveiled a charged, rotating black hole solution within the framework of low-energy heterotic string theory, termed the Kerr-Sen black hole. This solution modifies the conventional action of general relativity by incorporating supplementary fields derived from heterotic string theory, as delineated by[94, 99],
@@
$\begin{split}S=\int d^{4}x\sqrt{-\widetilde{g}}e^{-\Phi}\left[\mathcal{R}+(\nabla\Phi)^{2}-\frac{1}{8}F^{2}-\frac{1}{12}H^{2}\right],\end{split}$ (4.1)
In the given context, $\tilde{g}$ represents the determinant of the metric tensor, denoted as $g_{\mu\nu}$. The Ricci scalar is symbolized by $\mathcal{R}$, which is a key scalar quantity in General Relativity that represents curvature. The term $F$ is defined as $F_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}$ is the field strength tensor associated with the $U(1)$ Maxwell field, encapsulating the electromagnetic field’s properties. The scalar dilaton field is represented by $\Phi$, which arises in theories of gravity and string theory as a means of introducing a varying gravitational constant. Lastly, $H$ is expressed as $H_{\mu\nu\rho}H^{\mu\nu\rho}$, indicating the field strength of the axion field, a hypothetical elementary particle postulated to account for the CP problem in quantum chromodynamics. A conformal transformation of the metric is employed to transition to the Einstein frame, which is a particular representation of the gravitational field equations where the tensorial part of the action takes the simplest form. This transformation adjusts the metric tensor while preserving the angle between vectors, thereby simplifying the action’s structure in relation to the curvature of spacetime[94, 99],
$\begin{split}ds^{2}_{E}=e^{-\Phi}d\widetilde{s}^{2}.\end{split}$ (4.2)
So, we will have $S$ in the Einstein frame as follows[94, 99],
$\begin{split}S=\int d^{4}x\sqrt{-g}\left[R-\frac{1}{2}(\nabla\Phi)^{2}-\frac{e^{-\Phi}}{8}F^{2}-\frac{e^{-2\Phi}}{12}H^{2}\right].\end{split}$ (4.3)
In the absence of dilaton, vector, and axion fields, the Einstein-Hilbert action is a specific case. To derive black holes with a nonzero cosmological constant, the three-form field $(H\equiv d\mathcal{B}-A\wedge\frac{F}{4}=-e^{2\phi}\star d\chi)$ can be reformulated, where $\mathcal{B}$ is an anti-symmetric two-form potential and the ()($\star$)( ⋆ ) operator represents the Hodge duality. As a result, the Lagrangian can be written in an alternative but equivalent form[93, 94],
$\hat{L}=\sqrt{-g}\left[R-\frac{1}{2}(\partial\phi)^{2}-\frac{1}{2}e^{2\phi}(\partial\chi)^{2}-e^{-\phi}F^{2}\right]+\frac{\chi}{2}\epsilon^{\mu\nu\rho\lambda}F_{\mu\nu}F_{\rho\lambda}+\sqrt{-g}\left[4+e^{-\phi}+e^{\phi}(1+\chi^{2})\right]/\ell^{2},$
where $(\epsilon^{\mu\nu\rho\lambda})$ is the four-dimensional Levi-Civita antisymmetric tensor density, and $(\ell)$ denotes the cosmological scale or the reciprocal of the gauge coupling constant. In Boyer-Lindquist coordinates, the metric for an AKS black hole is described by the following expression[94, 99],
$\begin{split}ds^{2}=-\frac{\Delta_{r}}{\rho^{2}}(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}[-\frac{(r^{2}+2br+a^{2})}{\Xi}d\phi+adt]^{2},\end{split}$ (4.4)
@@
$\begin{split}&\rho^{2}=r^{2}+a^{2}\cos^{2}\theta+2br,\\ &\Delta=(r^{2}+2br+a^{2})(1+\frac{r^{2}+2br}{\ell^{2}})-2mr,\\ &\Xi=1-\frac{a^{2}}{\ell^{2}},\\ &\Delta_{\theta}=1-\frac{a^{2}}{\ell^{2}}\cos^{2}\theta.\end{split}$ (4.5)
The $b$ parameter signifies the dyonic charge of black holes ($b=\widehat{q}^{2}/(2m)$), linked to their electric charge ($\widehat{q}$) and mass (m). $a$ is the rotation parameter. The Kerr-Sen black holes emerge in certain limits, while the GMGHS solution describes non-rotating cases. Pioneering work by Gibbons, Maeda, Garfinkle, Horowitz, and Strominger has led to a deeper understanding of these black holes, especially in the context of AdS spacetimes, where their mass, angular momentum, and charge are interrelated. Studies have focused on their shadows and thermodynamics, contributing to the broader field of black hole physics[105, 106, 107, 108, 109]. The $M$, $J$, and $Q$ are given by,
$\begin{split}&M=\frac{m}{\Xi^{2}},\\ &J=\frac{ma}{\Xi^{2}},\\ &Q=\frac{\widehat{q}}{\Xi}.\end{split}$ (4.6)
Also, we can obtain the entropy, mass, and temperature in terms of $r_{+}$ and $a$, $b$ as follows[94],
$\begin{split}S=\frac{A}{4}=\frac{\pi(r_{+}^{2}+2br+a^{2})}{\Xi},\end{split}$ (4.7)
$\begin{split}M=\frac{\left(a^{2}+2br+r^{2}\right)\left(\frac{2br+r^{2}}{l^{2}}+1\right)}{2\Xi^{2}r},\end{split}$
$\begin{split}T=\frac{a^{2}\left(r^{2}-l^{2}\right)+r^{2}\left(4b^{2}+8br+l^{2}+3r^{2}\right)}{4\pi l^{2}r\left(a^{2}+2br+r\right)}\end{split}$
In the context of black hole physics, the variables $r_{+}$, $M$, $J$, and $Q$ represent the event horizon radius, mass, angular momentum, and charge, respectively. The event horizon radius $r_{+}$ is specifically the largest solution to the equation $\Delta=0$. Setting the parameter b to zero aligns the thermodynamic properties with those of the Kerr-AdS black holes, which are characterized as rotating solutions to the four-dimensional Einstein field equations with a negative cosmological constant. Furthermore, the critical temperature and pressure for the AKS black hole have been determined through specific calculations as follows[109],
$\begin{split}&T_{c}=-192J^{2}Q^{2}\bigg{(}\pi(26v^{2}+14v-33)-27\bigg{)}v^{3}\\ &+\bigg{(}v^{4}\bigg{[}36864J^{4}\bigg{(}153Q^{4}v^{2}+\pi^{2}Q^{4}v^{2}\big{(}26v^{2}+14v-33\big{)}^{2}-6\pi Q^{4}v^{2}\big{(}90v^{2}+206v-425\big{)}+72\pi Q^{2}\bigg{)}\\ &+384\pi J^{2}Q^{2}v^{4}\bigg{(}\pi\big{(}12Q^{2}-1\big{)}\big{(}26v^{2}+14v-33\big{)}-228Q^{2}+3\bigg{)}+\pi^{2}(1-12Q^{2})^{2}v^{6}\bigg{]}\bigg{)}^{1/2}\\ &+12\pi Q^{2}v^{5}-\pi v^{5}\bigg{/}\bigg{(}4608\pi J^{2}Q^{2}v^{4}\bigg{)},\end{split}$
and
$\begin{split}&P_{c}=\frac{(9+11\pi)\bigg{[}J^{2}Q^{2}(12Q^{2}-1)\bigg{(}-4(19+33\pi)Q^{2}+11\pi+1\bigg{)}v^{9}\bigg{(}768J^{2}Q^{2}(13v+7)+(12Q^{2}-1)v\bigg{)}\bigg{]}^{1/2}}{\sqrt{2}\pi v^{7}}\\ &+\bigg{[}12(416J^{2}+1)Q^{2}-1\bigg{]}\bigg{(}v^{8}\pi(12Q^{2}-1)v\bigg{[}768J^{2}Q^{2}(13v+7)+12Q^{2}v-1\bigg{]}\\ &+1152J^{2}\bigg{(}-4(19+33\pi)Q^{2}+11\pi+1\bigg{)}Q^{2}\bigg{)}^{1/2}\bigg{/}\bigg{(}9216\sqrt{\pi}J^{2}Q^{2}v^{5}\bigg{)}\\ &+\frac{(12Q^{2}-1)\big{(}12(832J^{2}+1)Q^{2}-1\big{)}}{9216J^{2}Q^{2}}-\frac{(9+11\pi)(12Q^{2}-1)}{24\pi v^{2}}+\frac{7(12Q^{2}-1)}{24v}.\end{split}$
where $v$ is the new variable which relates to the specific volume of the van der Waals fluid as[109],
$\begin{split}v=2\big{(}\frac{3V}{4\pi}\big{)}^{1/3}=2r_{+}+2b+\frac{12J^{2}}{\big{(}8\pi Pr_{+}^{4}+3r_{+}^{2}\big{)}^{2}}\big{(}4b+\frac{80}{3}Pbr_{+}^{2}+8\pi Pr_{+}^{3}+3r_{+}\big{)}.\end{split}$
@@ 5 Discussion and results
In this section, we investigate the JTE process for AdS black holes: AEMPYM and AKS and derive a clear formula for the JTC. We also examine how the charge and the parameters of AEMPYM and Kerr-Sen theories affect the inversion curves. We also compare the inversion curves for different scenarios
@@ 5.1 Case I
In this section, we explore the JTE of black hole systems in the extended phase space, where the black hole mass M is the same as the enthalpy H. In our study, we conduct a short comparison of our results with the (JT) process as applied to other black holes. These comparative results are summarized in table 5. This allows us to evaluate the consistency and discrepancies between our findings and the other data within this concept. Using [27, 71] and considering the mass of black holes, we can write the pressure P as a function of M and $r_{+}$, the mass and the horizon radius of a black hole. By substituting this expression in the temperature formula, we can obtain the temperature as a function of M and $r_{+}$ as well. Furthermore, we can express the mass M and the temperature T in terms of the pressure P and the radius $r_{+}$ of a black hole. Therefore, we get,
$P(M,r_{+})=-\frac{3\left(\frac{2C^{2}(n-1)r_{+}^{-2n}}{n}+Qr_{+}^{-4q}+\frac{1}{r_{+}^{2}}\right)}{8\pi}+\frac{6Mr_{+}^{-n-1}}{n\omega_{n}}.$ (5.1)
@@
$T(M,r_{+})=-\frac{2C^{2}(n-1)r_{+}^{2-2n}+2qQr_{+}^{2-4q}+1}{2\pi r_{+}}+\frac{4M(n+1)r_{+}^{-n}}{n\omega_{n}}.$ (5.2)
Due to the complexity of the aforementioned such as $T(M,P)$ equations, an analytical solution is not feasible. Therefore, we have resorted to numerical solutions to further our work. A portion of these calculations can be found in the appendix. According to the definition of the JTC $\mu=\left(\frac{\partial T}{\partial P}\right)_{M}$, the inversion pressure and temperature between the cooling and heating regions are ($\tilde{P}$-$\tilde{T}$), which are determined by $\mu=0$. Therefore, the most important thing is to find the function expression of $\mu$. By setting $\mu=0$, one can obtain the inversion points ($\tilde{P}$, $\tilde{T}$) for different fixed enthalpy M. With respect to[23], the JTC is given by,
$\begin{split}\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{M}=\frac{1}{C_{P}}\big{[}T(\frac{\partial V}{\partial T}\big{)}_{P}-V\big{]}.\end{split}$ (5.3)
@@
$\begin{split}\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{M}=\frac{\partial_{,r_{+}}T}{\partial_{,r_{+}}P}.\end{split}$ (5.4)
@@
$\begin{split}&\mu=\frac{2nr_{+}\bigg{(}\omega_{n}\bigg{[}r_{+}^{4q}\big{(}2C^{2}(2n^{2}-3n+1)r_{+}^{2}+r_{+}^{2n}\big{)}+8q^{2}Qr^{2n+2}-2qQr_{+}^{2n+2}\bigg{]}-8\pi M(n+1)r_{+}^{n+4q+1}\bigg{)}}{3\bigg{(}n\omega_{n}\bigg{[}r_{+}^{4q}\big{(}2C^{2}(n-1)r_{+}^{2}+r_{+}^{2n}\big{)}+2qQr_{+}^{2n+2}\bigg{]}-8\pi M(n+1)r^{n+4q+1}\bigg{)}}.\end{split}$ (5.5)
@@
$T_{i}=V\big{(}\frac{\partial T}{\partial V}\big{)}_{P},$ (5.6)
$\begin{split}V=\frac{\partial M}{\partial P}=\frac{1}{6}nr_{+}^{n+1}\omega_{n}.\end{split}$ (5.7)
@@
$\begin{split}T_{i}=V\frac{\partial T}{\partial V}=V\frac{\frac{\partial T}{\partial r_{+}}}{\frac{\partial V}{\partial r_{+}}}.\end{split}$ (5.8)
@@
This is beneficial for identifying the areas of heating and cooling in the $T-P$ plane. We calculate $T_{i}$ using the Eqs. (5.2), (5.5), (5.6), (5.7) and (5.8),
$\begin{split}T_{i}=\frac{r_{+}\left(\frac{6C^{2}(n-1)^{2}(2n-1)r_{+}^{-2n}}{n}+8\pi(n+1)P_{i}+3(1-4q)Q(n-4q+1)r_{+}^{-4q}+\frac{3-3n}{r_{+}^{2}}\right)}{12\pi(n+1)}.\end{split}$ (5.9)
We can also have from Eq. (5.2),
$\begin{split}T_{i}=\frac{-\frac{6C^{2}(n-1)^{2}r_{+}^{1-2n}}{n}+8\pi(n+1)r_{+}P_{i}+3Q(n-4q+1)r_{+}^{1-4q}+\frac{3(n-1)}{r_{+}}}{12\pi}.\end{split}$ (5.10)
Based on the above equation, we can derive an expression commonly denoted as $(T^{min}_{i})$. This is achieved by setting the value of $(P_{i})$ to zero. Also, we can rewrite the Eqs. (5.9) and (5.11) with respect to Eq. (3.5) as follows,
$\begin{split}T^{min}_{i}=\frac{r_{+}^{-2n-4q-1}\left(2C^{2}(n-1)^{2}(2n-1)r_{+}^{4q+2}+(4q-1)r_{+}^{2n+2}\left((n-1)nQ_{1}^{2}\right){}^{q}-(n-1)nr_{+}^{2(n+2q)}\right)}{4\pi n(n+1)},\end{split}$
and
$\begin{split}T^{min}_{i}=\frac{r_{+}^{-2n-4q-1}\left(-2C^{2}(n-1)^{2}r_{+}^{4q+2}+r_{+}^{2n+2}\left(-\left((n-1)nQ_{1}^{2}\right){}^{q}\right)+(n-1)nr_{+}^{2(n+2q)}\right)}{4\pi n}.\end{split}$
@@
Fig. (1) displays the Hawking temperature as a function of the horizon. We keep the free parameters constant for each plot. In each subfigure, we can observe some zero points for different free parameters. These zero points correspond to the divergence points of the JTC, as we can easily see in Fig. (1). According to the radius of the horizon and the values of free parameters of a black hole, for small values for the radius of the event horizon, our structural behavior is completely distinct, and for larger radii, the figures almost converge. The differences between our research and the referenced study are discernible. For instance, while our work presents findings in a specific context, the [65] may approach the subject from a different angle. By carefully referencing the other study, we not only acknowledge its contributions but also reinforce the originality and integrity of our research
Figure 1: The plot of temperature $T$ in terms of $r_{+}$ at $P=0.5,C=1/4,q=1/3$, and $Q=0.5$ for different values of $n$ in Fig. (1), at $P=0.5,n=2,C=0.5$, and $Q=0.5$ for different values of $q$ in Fig. (1), at $P=0.5,n=2,q=0.5$, and $Q=0.5$ for different values of $C$ in Fig. (1) and at $P=0.5,n=2,q=0.5$, and $C=0.05$ for different values of $Q$ in Fig. (1) for the AEMPYM black hole
@@
We have plotted the isenthalpic curves and the inversion curve of the AEMPYM black hole for various values of the free parameter in each plot of Fig. (2). In every subfigure, two isenthalpic curves with different values of $M$, along with the corresponding inversion curve occur at the highest point of the isenthalpic curves. We denote the inversion temperature and pressure of each isenthalpic curve as $T_{i}$ and $P_{i}$, respectively. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole undergoes cooling during the expansion process. However, for $P>P_{i}$, the isenthalpic curve has a negative slope, implying that the black hole experiences heating during the expansion process. This behavior is consistent for different values of the free parameter in Figs. ((2)-(2)).
Figure 2: The plot of isenthalpic and inversion curves for different values of $M$. In Fig. (2) at $n=2,C=1/4,Q=0.5$, and $q=1/3$, in Fig. (2) at $n=3,q=1.5,C=0.5$, and $Q=0.5$ for the AEMPYM black hole
@@
However, we can illustrate the relation between the inversion temperature $T_{i}$ and the inversion pressure $P_{i}$ by using the equations given above. Fig. (3) shows the inversion curves for different free parameters. The inversion curve has only one branch. The inversion temperature rises steadily with the inversion pressure, but the slope of the inversion curves becomes smaller. We can also notice some fine structures in the cases of subfigures (3). For low pressure, the inversion temperature varies with the free parameters that are specified in each plot. It was interesting for us to study the effect of parameters and we observed that the slope of the inversion curve increases with the changes in the various values of each of the parameters.
Figure 3: The plot of inversion curves ($T_{i}-P_{i}$). In Fig. (3) at $n=2,C=0.5$, and $Q=1$ for different values of $q$, in Fig. (3) at $q=1.5,C=0.5$, and $Q=1$ for different values of $n$, in Fig. (3) at $n=2,q=1.5$, and $Q=1$ for different values of $C$ and in Fig. (3) at $n=2,q=1.5$, and $C=0.5$ for different values of $Q$ for the AEMPYM black hole